It has been reported that 2% of all cars on the highway are traveling at speeds more than 130 km/h. If the speeds of six random cars are measured via radar, what is the probability that at least one car is going more than 130 km/h? 


  • 0.114 

  • 0.886 

  • 0.226 

  • 0.774 
Answer #1

P(speed of a single car is more than 130km/h)=0.02

n=6

P(atleast one car is going more than 130km/h)=1-P(no car is going more than 130km/h)

\\ =1-(1-0.02)^6\\ \\ =0.1141576\\ \\ =0.114